# ChangeMamba: Remote Sensing Change Detection with Spatio-Temporal State Space Model

### v2

@@ Abstract
Convolutional neural networks (CNN) and Transformers have made impressive progress in the field of remote sensing change detection (CD). However, both architectures have inherent shortcomings. Recently, the Mamba architecture, based on state space models, has shown remarkable performance in a series of natural language processing tasks, which can effectively compensate for the shortcomings of the above two architectures. In this paper, we explore for the first time the potential of the Mamba architecture for remote sensing CD tasks. We tailor the corresponding frameworks, called MambaBCD, MambaSCD, and MambaBDA, for binary change detection (BCD), semantic change detection (SCD), and building damage assessment (BDA), respectively. All three frameworks adopt the cutting-edge Visual Mamba architecture as the encoder, which allows full learning of global spatial contextual information from the input images. For the change decoder, which is available in all three architectures, we propose three spatio-temporal relationship modeling mechanisms, which can be naturally combined with the Mamba architecture and fully utilize its attribute to achieve spatio-temporal interaction of multi-temporal features, thereby obtaining accurate change information. On five benchmark datasets, our proposed frameworks outperform current CNN- and Transformer-based approaches without using any complex training strategies or tricks, fully demonstrating the potential of the Mamba architecture in CD tasks. Specifically, we obtained 83.11, 88.39 and 94.19 F1 scores on the three BCD datasets SYSU, LEVIR-CD+, and WHU-CD; on the SCD dataset SECOND, we obtained 24.11 SeK; and on the BDA dataset xBD, we obtained 81.41 overall F1 score. Further experiments show that our architecture is quite robust to degraded data. The source code is available in https://github.com/ChenHongruixuan/MambaCD.
@@
The emergence of deep learning has brought new models and paradigms for CD, greatly improving the efficiency and accuracy of CD. After the early days of developing image patch-based approaches with tiny deep learning models [16, 17, 18], the released of a number of large-scale benchmark datasets [19, 20, 21] have allowed us to start designing and introducing larger and deeper models. Convolutional neural network (CNN)-based methods have been dominant within the field of CD since the introduction of fully convolutional networks (FCNs) to the field by Daudt et al. [22]. By employing models within the computer vision field as well as a priori knowledge of the CD task, many representative methods have been proposed [23, 24, 3, 4, 25, 26, 27]. Despite the decent results achieved by these methods, the inherent shortcomings of the CNN model structure, i.e., the inability of the limited receptive field to capture long-range dependencies between pixels, make these methods still fall short when dealing with complex and diverse multi-temporal scenes in images with different spatial-temporal resolutions [28]. The emergence of visual Transformer [29] provides an effective way to solve the above problems. By means of the stacked self-attention modules, the Transformer architecture can efficiently model the relationship between all the pixels of the whole image. Currently, more and more CD architectures adopt Transformer architectures as the encoder to extract representative and robust features [30, 31, 32] and utilize it in the decoder to capture the spatio-temporal relationships between multi-temporal features [28, 33], exhibiting superior performance.
@@
Nonetheless, the self-attention operation demands quadratic complexity in terms of image sizes, resulting in expensive computational overhead, which is detrimental in dense prediction tasks, e.g., land-cover mapping, object detection, and CD, in large-scale remote sensing datasets. Some available solutions, such as limiting the size of the computational window or step size to improve the efficiency of attention, are at the cost of imposing limits on the size of the reception field [34, 35]. As a viable alternative to the Transformer architecture, state space sequence models (SSMs), especially structured state space sequence models (S4) [36], have shown cutting-edge performance in continuous long-sequence data analysis, with the promising attribute of linearly scaling in sequence length. The Mamba architecture [37] further improves on S4 by employing a selection mechanism that allows the model to select relevant information in an input-dependent manner. By combining it with hardware-aware implementations, Mamba outperforms Transformers on a number of downstream tasks with high efficiency. Very recently, Mamba architecture has been extended to image data and shown promising results on some visual tasks [38, 39]. However, its potential for high-resolution optical images and different CD subtasks is still under-explored. How to design a suitable architecture based on a priori knowledge of CD tasks is undoubtedly valuable for subsequent research in the community.
@@
In this paper, we present the first attempt at introducing Mamba architecture for change detection tasks and proposed several spatio-temporal state space models (collectively called ChangeMamba) to realize effective and robust CD. ChangeMamba can effectively model the global spatial context and spatio-temporal relationships in multitemporal images, thereby showing very promising results on the three subtasks of CD, namely BCD, MCD, and BDA. Specifically, ChangeMamba is based on the recently proposed VMamba architecture [38], which adopts a Cross-Scan Module (CSM) to unfold image patches in different spatial directions to achieve effective modeling of global contextual information of images. Since CD tasks require the detector to adequately learn spatio-temporal features from multi-temporal images, we design three spatio-tepmoral relationship modelling mechanisms and naturally combine it with Mamba architecture, which can adequately model the spatio-temporal relationships between multi-temporal features by scanning them in different ways, and thus efficiently detect the different categories of changes, including binary changes, semantic changes, and building damage levels.
@@
Based on the Mamba architecture and incorporating the unique characteristics of CD tasks, we design the corresponding network frameworks for each of the three CD tasks: BCD, SCD, and BDA.
@@ II Related Works
In this section, we focus on reviewing deep learning-based CD methods concentrating on optical high-resolution remote sensing images as data sources.
@@ II-A CNN-Based Method
The rapid evolution of CNN and their prowess in automatically extracting hierarchical features have led to wide applications in CD tasks. After the early stage of patch-based framework [16, 18], Daudt et al. [22] presented the first attempt at introducing FCNs into the BCD field and three basic FCNs architecture were developed. After that, many representative network architectures were proposed. Zhang et al. [23] sought to tailor the networks for semantic segmentation task more suitable for the BCD task by integrating feature change maps as a deep supervised signal, thus refining the final prediction. To make the framework suitable for handling multi-source images, Chen et al. [17] proposed a framework by combining long- and short-term memory (LSTM) networks with CNNs.Furthermore, Fang et al. [40] designed a Siamese architecture with denser connections to facilitate more effective shallow information exchange for BCD. To combat the challenge of uneven foreground-background distributions within BCD, Han et al. [41] employed a novel sampling approach and an attention mechanism to integrate global information more effectively. After that, they introduced an innovative self-attention module that not only enhances long-distance dependency, but also specifically targets the refinement of BCD predictions, significantly improving the delineation of change area boundaries and reducing the occurrence of inner holes [42]. Zhao et al. [43] explored the effect of different fusion strategies on BCD performance and designed a dual encoder-decoder architecture to enhance the detection performance. To solve the problem of insufficient sample, the transfer learning technique was also introduced. Guo et al. [4] developed an attention-based network for semi-supervised building updating. Cao et al. [25] designed a full-level fused cross-task transfer learning architecture for building change detection. Recently, Chen et al. [44] has developed a sample generation method based on simply exchanging image patches, which can train change detectors on single-temporal images in unsupervised manner.
@@
Compared to BCD, which only aims to identify “where” changes have occurred, SCD further seeks to determine “what” the changes are. Although this poses a greater challenge, it is more important for practical application analysis [45, 24]. Early explorations based on CNNs for SCD is the study [45], which adopts a multitask learning framework to simultaneously predict binary change maps and land cover maps, using the latter to assist in generating the final semantic change prediction. Mou et al. [46] proposed to combine recurrent neural network (RNN) and CNN for SCD, where a Siamese CNN extracts semantic features and RNN models temporal dependencies to classify multiple types of changes. Yang et al. [20] introduced a benchmark dataset for SCD along with a new evaluation metric. Based on the framework in [45], Ding et al. [47] proposed a more powerful architecture to address the insufficient exchange of information between the semantic and change decoders. A new loss functions was introduced to force the network to correlate two decoders. Zheng et al. [24] continues to strengthen the interactions between multi-temporal features, designing a temporal-symmetric transformer (TST) module to model the spatio-temporal relationships. In [48], channel attention was used to embed change information into temporal features. In [49], temporal salient features were fused to extract and improve the representation of changes. In [50] and [14], unsupervised SCD methods were also studied, using pre-trained CNNs or unsupervised learning techniques to extract hierarchical features, then comparing these features to determine changes.
@@
Due to lack of suitable benchmark datasets caused by the scarcity of disaster events and difficulty in annotation, the methodology progression of BDA has been relatively slow. The establishment of xBD benchmark dataset of the xView2 Challenge [51], in turn, provides a large-scale, common benchmark for the field. The procedural approach to BDA typically encompasses two primary phases: identifying the location of the building and assessing the level of damage incurred. In the xBD dataset, there are four damahge levels, i.e., “no damage”, “minor damage”, “major damage”, and “destroyed”. This bifurcation essentially positions BDA as a nuanced subset of SCD, with a focus on quantifying damage levels. Among established approaches, the xView2 challenge [51] introduced a baseline method that employs a residual UNet for building localization and damage classification, trained with pre and post-disaster imagery. The Siamese-UNet, the winning approach of the xView2 challenge [51], utilized a UNet framework for both pixel-level building localization and damage level classification. It initialized the UNet for the damage classification task with the parameters from the UNet for the building localization task. Further advancements [52] utilized Mask-RCNN architecture [53] to address BDA, incorporating Siamese encoders, multitask heads, and instance-level predictions. Zheng et al. [3] introduced a deep object-based architecture that encompasses shared and task-specific encoders, alongside an FCN-based multitask decoder. This development enhances task integration with an object-based post-processing strategy, resulting in more refined outcomes.
@@
The development of CNN-based CD methods has matured significantly over time. However, despite their excellent ability to extract local features, CNN architectures inherently struggle to capture long-distance dependencies [29]. This limitation becomes particularly critical in CD tasks, where change regions, compared to background areas, are often sparse and scattered. The network’s ability to model relationships beyond local areas is essential for accurately identifying these changed regions [28, 30].
@@ II-B Transformer-Based Method
Recently, with the promising results obtained by Transformers in the computer vision fields [29, 35], their superior long-distance modeling capabilities, have made them effectively overcome the challenges faced by the CNNs. As a result, many studies have begun to address CD tasks using Transformer architectures. In the field of BCD, Chen et al. [28] emerged as a pioneering approach, transforming multi-temporal images into semantic tokens to model spatial-temporal relationships within a token-based framework, enhancing the change detection process. Bandara et al. [30] proposed a pure transformer-based Siamese network, which has a Siamese Transformer encoder coupled with an multilayer perceptron (MLP) decoder. This architecture eliminates the need for a CNN-based feature extractor. Similarly, Zhang et al. [32] proposed a pure Transformer-driven model, adopting a Siamese U-shaped structure to fully recover detailed change information. Specifically, it leveraged the Swin Transformer blocks [34] across its architecture, further demonstrating the transformer’s capability in BCD tasks. Li et al. [33] introduced a hybrid model that blends the global contextual modelling capability of Transformers with the CNN block and UNet architecture. It eliminates UNet’s redundant information and enhances feature quality through a difference enhancement module, leading to more precise detection results.
@@
Transformer-based BDA methods are currently understudied. Chen et al. [31] proposed the first pure Transformer-based BDA architecture called DamFormer. It has a Siamese Mix Transformer (MiT) [35] encoder and lightweight all-MLP decoder to efficiently process multi-temporal image pairs. DamFormer employs a multi-temporal fusion module for improved information integration, and a dual-task decoder for precise building localization and damage classification. By adopting strategies like weight sharing and innovative fusion mechanisms, it efficiently handles high-resolution imagery with lower computational costs, making it effective and streamlined for BDA.
Although Transformer-based methods have significantly enhanced CD accuracy through their powerful nonlocal modeling capabilities, they come with substantial computational resource demands. Specifically, as the number of tokens increases, the consumption of computational resources escalates quadratically, posing challenges for processing large-scale input, such as optical high-resolution imagery. Recently, the emerging Mamba architecture not only rivals Transformers in non-local modeling capabilities but also maintains computational overhead at a linear growth rate, which opens up the possibility of efficient and accurate CD.
@@ II-C State Space Model
The SSM concept first popped up with the S4 model [36], offering a new way to handle contextual information globally, which caught attention because of its attractive property of scaling linearly in sequence length. Based on S4, Smith et al. [56] proposed a new S5 model by introducing MIMO SSM and efficient parallel scan into the S4 model. After that, the H3 model [57] further advanced these concepts, achieving performance on par with Transformers in language modeling tasks. Recently, Gu et al. [37] proposed a data-dependent SSM layer and built a generic language model backbone called Mamba, which outperforms Transformers at different scales on large-scale real data and scales linearly in sequence length. Very recently, VMamba [38] and Vision Mamba [39] have extended Mamba architecture to 2D image data, showing superior performance on many computer vision tasks. Inspired by this progress, some pioneering work has introduced the Mamba architecture to the field of remote sensing image processing [58, 59]. Their studies show that the Mamba architecture can yield performance comparable to the advanced CNN and Transformer architectures on scene classification [59] and pansharpening [58] tasks. However, these works still focus on either low-level tasks or classification tasks, and all of these tasks are single-temporal tasks [44]. The potential of the Mamba architecture for multi-temporal remote sensing image-related scenarios and intensive prediction tasks remains to be explored.

### v3

@@ Abstract
Convolutional neural networks (CNN) and Transformers have made impressive progress in the field of remote sensing change detection (CD). However, both architectures have inherent shortcomings: CNN are constrained by a limited receptive field that may hinder their ability to capture broader spatial contexts, while Transformers are computationally intensive, making them costly to train and deploy on large datasets. Recently, the Mamba architecture, based on state space models, has shown remarkable performance in a series of natural language processing tasks, which can effectively compensate for the shortcomings of the above two architectures. In this paper, we explore for the first time the potential of the Mamba architecture for remote sensing CD tasks. We tailor the corresponding frameworks, called MambaBCD, MambaSCD, and MambaBDA, for binary change detection (BCD), semantic change detection (SCD), and building damage assessment (BDA), respectively. All three frameworks adopt the cutting-edge Visual Mamba architecture as the encoder, which allows full learning of global spatial contextual information from the input images. For the change decoder, which is available in all three architectures, we propose three spatio-temporal relationship modeling mechanisms, which can be naturally combined with the Mamba architecture and fully utilize its attribute to achieve spatio-temporal interaction of multi-temporal features, thereby obtaining accurate change information. On five benchmark datasets, our proposed frameworks outperform current CNN- and Transformer-based approaches without using any complex training strategies or tricks, fully demonstrating the potential of the Mamba architecture in CD tasks. Specifically, we obtained 83.11, 88.39 and 94.19 F1 scores on the three BCD datasets SYSU, LEVIR-CD+, and WHU-CD; on the SCD dataset SECOND, we obtained 24.11 SeK; and on the BDA dataset xBD, we obtained 81.41 overall F1 score. Further experiments show that our architecture is quite robust to degraded data. The source code is available in https://github.com/ChenHongruixuan/MambaCD.
@@
The emergence of deep learning has brought new models and paradigms for CD, greatly improving the efficiency and accuracy of CD. After the early days of developing image patch-based approaches with tiny deep learning models [16, 17, 18], the release of a number of large-scale benchmark datasets [19, 20, 21] has allowed us to start designing and introducing larger and deeper models. Convolutional neural network (CNN)-based methods have been dominant within the field of CD since the introduction of fully convolutional networks (FCNs) to the field by Daudt et al. [22]. By employing models within the computer vision field as well as a priori knowledge of the CD task, many representative methods have been proposed [23, 24, 3, 4, 25, 26, 27]. Despite the decent results achieved by these methods, the inherent shortcomings of the CNN model structure, i.e., the inability of the limited receptive field to capture long-range dependencies between pixels, make these methods still fall short when dealing with complex and diverse multi-temporal scenes in images with different spatial-temporal resolutions [28]. The emergence of the visual Transformer [29] provides an effective way to solve the above problems. By means of the stacked self-attention modules, the Transformer architecture can fully model the relationship between all the pixels of the whole image. Currently, more and more CD architectures adopt Transformer architectures as the encoder to extract representative and robust features [30, 31, 32] and utilize it in the decoder to capture the spatio-temporal relationships between multi-temporal features [28, 33], exhibiting superior performance.
@@
Nonetheless, the self-attention operation demands quadratic complexity in terms of image sizes, resulting in expensive computational overhead, which is detrimental in dense prediction tasks, e.g., land-cover mapping, object detection, and CD, in large-scale remote sensing datasets. Some available solutions, such as limiting the size of the computational window or the step size to improve attention efficiency, come at the cost of imposing limits on the size of the reception field [34, 35]. As a viable alternative to the Transformer architecture, state space models (SSMs), especially structured state space sequence models (S4) [36], have shown cutting-edge performance in continuous long-sequence data analysis, with the promising attribute of linear scaling in sequence length. The Mamba architecture [37] further improves on the S4 model by employing a selection mechanism that allows the model to select relevant information in an input-dependent manner. By combining it with hardware-aware implementations, Mamba outperforms Transformers on a number of downstream tasks with high efficiency. Very recently, Mamba architecture has been extended to image data and shown promising results on some visual tasks [38, 39]. However, its potential for high-resolution optical images and different CD subtasks is still under-explored. How to design a suitable architecture based on a priori knowledge of CD tasks is undoubtedly valuable for subsequent research in the community.
@@
In this paper, we present the first attempt at introducing Mamba architecture for CD tasks and propose several spatio-temporal state space models (collectively called ChangeMamba) to realize effective and robust CD. ChangeMamba can effectively model the global spatial context and spatio-temporal relationships in multi-temporal images, thereby showing very promising results on the three subtasks of CD, namely BCD, MCD, and BDA. Specifically, ChangeMamba is based on the recently proposed VMamba architecture [38], which adopts a Cross-Scan Module (CSM) to unfold image patches in different spatial directions to achieve effective modeling of global contextual information from images. Since CD tasks require the detector to adequately learn spatio-temporal features from multi-temporal images, we design three spatio-tepmoral relationship modeling mechanisms and naturally combine them with Mamba architecture. The proposed three mechanisms can adequately model the spatio-temporal relationships between multi-temporal features by scanning them in different ways, and thus effectively detect the different categories of changes, including binary changes, semantic changes, and building damage levels.
@@
Based on the Mamba architecture and incorporating the unique characteristics of CD tasks, we design the corresponding network frameworks for each of the three CD tasks, i.e., BCD, SCD, and BDA.
@@ II Related Works
In this section, we focus on reviewing deep learning-based CD methods using optical high-resolution remote sensing images as data sources.
@@ II-A CNN-Based Method
The rapid evolution of CNN and its prowess in automatically extracting hierarchical features have led to wide applications in CD tasks. After the early stage of the patch-based framework [16, 18], Daudt et al. [22] presented the first attempt to introduce FCNs into the BCD field and three basic FCNs architectures were developed. After that, many representative network architectures were proposed. Zhang et al. [23] sought to tailor networks for semantic segmentation tasks more suitable for the BCD task by integrating feature change maps as a deep supervised signal, thus refining the final prediction. To make the framework suitable for handling multi-source images, Chen et al. [17] proposed a framework by combining long- and short-term memory (LSTM) networks with CNNs. Furthermore, Fang et al. [40] designed a Siamese architecture with denser connections to facilitate more effective shallow information exchange for BCD. To combat the challenge of uneven foreground-background distributions within BCD, Han et al. [41] employed a novel sampling approach and an attention mechanism to integrate global information more effectively. After that, they introduced an innovative self-attention module that not only enhances long-distance dependency, but also specifically targets the refinement of BCD predictions, significantly improving the delineation of change area boundaries and reducing the occurrence of inner holes [42]. Zhao et al. [43] explored the effect of different fusion strategies on BCD performance and designed a dual encoder-decoder architecture to improve detection performance. To solve the problem of insufficient samples, the transfer learning technique was also introduced. Guo et al. [4] developed an attention-based network for semi-supervised building updating. Cao et al. [25] designed a full-level fused cross-task transfer learning architecture for building change detection. Recently, Chen et al. [44] has developed a sample generation method by simply exchanging image patches, which can train change detectors on single-temporal images in an unsupervised manner.
@@
Compared to BCD, which only aims to identify “where” changes have occurred, SCD further seeks to determine “what” the changes are. Although this poses a greater challenge, it is more important for practical application analysis [45, 24]. Early explorations based on CNNs for SCD is the study [45], which adopts a multitask learning framework to simultaneously predict binary change maps and land cover maps, using the latter to assist in generating the final semantic change prediction. Mou et al. [46] proposed to combine recurrent neural network (RNN) and CNN for SCD, where a Siamese CNN extracts semantic features and RNN models temporal dependencies to classify multiple types of changes. Yang et al. [20] introduced a benchmark dataset for SCD along with a new evaluation metric. Based on the framework in [45], Ding et al. [47] proposed a more powerful architecture to address the insufficient exchange of information between the semantic and change decoders. A new loss function was introduced to force the network to correlate two decoders. Zheng et al. [24] continues to strengthen the interactions between multi-temporal features, designing a temporal-symmetric transformer (TST) module to model the spatio-temporal relationships. In [48], channel attention was used to embed change information into temporal features. In [49], temporal salient features were fused to extract and improve the representation of changes. In [50] and [14], unsupervised SCD methods were also studied, using pre-trained CNNs or unsupervised learning techniques to extract hierarchical features, then comparing these features to determine changes.
@@
Due to the lack of suitable benchmark datasets caused by the scarcity of disaster events and difficulty in annotation, the methodology progression of BDA has been relatively slow. The establishment of the xBD benchmark dataset in the xView2 Challenge [51], in turn, provides a large-scale, common benchmark for the field. The procedural approach to BDA typically encompasses two primary components: identifying the location of the building and assessing the level of damage incurred. In the xBD dataset, there are four damage levels, i.e., “no damage”, “minor damage”, “major damage”, and “destroyed”. This bifurcation essentially positions BDA as a nuanced subset of SCD, with a focus on quantifying damage levels. Among established approaches, the xView2 challenge [51] introduced a baseline method that employs a residual UNet for building localization and damage classification, trained with pre and post-disaster imagery. The Siamese-UNet, the winning approach of the xView2 challenge [51], utilized a UNet framework for both pixel-level building localization and damage-level classification. It initialized the UNet for the damage classification task with the parameters from the UNet for the building localization task. A further advancement [52] utilized Mask-RCNN architecture [53] to address BDA, incorporating Siamese encoders, multitask heads, and instance-level predictions. Zheng et al. [3] introduced a deep object-based architecture that encompasses shared and task-specific encoders, alongside an FCN-based multitask decoder. This development enhances task integration with an object-based post-processing strategy, resulting in more refined outcomes.
@@
The development of CNN-based CD methods has matured significantly over time. However, despite their excellent ability to extract local features, CNN architectures inherently struggle to capture long-distance dependencies [29]. This limitation becomes particularly critical in CD tasks, where changed regions, compared to background areas, are often sparse and scattered. The network’s ability to model relationships beyond local areas is essential for accurately identifying these changed regions [28, 30].
@@ II-B Transformer-Based Method
Recently, with the promising results obtained by Transformers in the computer vision field [29, 35], their superior long-distance modeling capabilities, have made them effectively overcome the challenges faced by the CNNs. As a result, many studies have begun to address CD tasks using Transformer architectures. In the field of BCD, Chen et al. [28] emerged as a pioneering approach, transforming multi-temporal images into semantic tokens to model spatial-temporal relationships within a token-based framework, enhancing the CD process. Bandara et al. [30] proposed a pure Transformer-based Siamese network, which has a Siamese Transformer encoder coupled with a multilayer perceptron (MLP) decoder. This architecture eliminates the need for a CNN-based feature extractor. Similarly, Zhang et al. [32] proposed a pure Transformer-driven model, adopting a Siamese U-shaped structure to fully recover detailed change information. Specifically, it leveraged the Swin Transformer blocks [34] across its architecture, further demonstrating the Transformer’s capability in BCD tasks. Li et al. [33] introduced a hybrid model that blends the global contextual modeling capability of Transformers with the CNN block and UNet architecture. It eliminates UNet’s redundant information and enhances feature quality through a difference enhancement module, leading to more precise detection results.
@@
Transformer-based BDA methods are currently understudied. Chen et al. [31] proposed the first pure Transformer-based BDA architecture called DamFormer. It has a Siamese Mix Transformer (MiT) [35] encoder and a lightweight all-MLP decoder to efficiently process multi-temporal image pairs. DamFormer employs a multi-temporal fusion module for improved information integration, and a dual-task decoder for precise building localization and damage classification. By adopting strategies like weight sharing and innovative fusion mechanisms, it efficiently handles high-resolution imagery with lower computational costs, making it effective and streamlined for BDA.
Although Transformer-based methods have significantly enhanced CD accuracy through their powerful nonlocal modeling capabilities, they come with substantial computational resource demands. Specifically, as the number of tokens increases, the consumption of computational resources escalates quadratically, posing challenges for processing large-scale input, such as optical high-resolution imagery. Recently, the emerging Mamba architecture [37] not only rivals Transformers in non-local modeling capabilities but also maintains computational overhead at a linear growth rate, which opens up the possibility of efficient and accurate CD.
@@ II-C State Space Model
The SSM concept first popped up with the S4 model [36], offering a new way to handle contextual information globally, which caught attention due to its attractive property of scaling linearly in sequence length. Based on the S4 model, Smith et al. [56] proposed a new S5 model by introducing MIMO SSM and efficient parallel scan into the S4 model. After that, the H3 model [57] further advanced these concepts, achieving a performance on par with that of Transformers in language modeling tasks. Recently, Gu et al. [37] proposed a data-dependent SSM layer and built a generic language model backbone called Mamba, which outperforms Transformers at different scales on large-scale real data and scales linearly in sequence length. Very recently, Visual Mamba [38] and Vision Mamba [39] have extended the Mamba architecture to 2D image data, showing superior performance on many computer vision tasks. Inspired by this progress, some pioneering work has introduced the Mamba architecture to the field of remote sensing image processing [58, 59]. Their studies show that the Mamba architecture can yield performance comparable to the advanced CNN and Transformer architectures on scene classification [59] and pansharpening [58] tasks. However, these works still focus on low-level tasks or classification tasks, and all of these tasks are single-temporal tasks [44]. The potential of the Mamba architecture for multi-temporal remote sensing image-related scenarios and dense prediction tasks remains to be explored.